A wire of resistance 5 O is drawn out so that its new length is 3 times its original length.What is the reistance of the new wire?
(A) 45 O (B) 15 O (C) 5/3 O (D) 5 O

Grade:

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

R = (rho)l/a

volume =l*a = constant

Now you can calculate

Regards
Vikas TU
3 years ago

By stretch the wire till its length is triple while the volume remains constant, the cross sectional area is three times less compared to its original value.

Therefore, when you computed the ratio between the original and the new value of the resistance, it would be:

R/R' = (L/A) / (L'/A') = 1/9
Thus, the new resistance would be R' = 9R = 45O